OpenStudy (anonymous):

Helpful steps please,Complete combustion of 2.50 g of a hydrocarbon produced 7.71 g of CO2 and 3.55 g of H2O. What is the empirical formula for the hydrocarbon? ( I tired a and I ended up with C2H4 and then I tried another way and got C2H23??)

OpenStudy (matt101):

Let's go through what you did and see if it makes sense. What was your first step?

OpenStudy (anonymous):

7.71g co2*1mol co2/44.01*1mol c/1 mol c= 0.175 3.55 g h20 1mol h20/18.01*2 H/1mol = 0.394 0.175 mol c/ 0.175= 1 C 0.394 mol h/0.175 = 2.3 H I got stuck around here